# Predictions for Detecting a Turndown in the Baryonic Tully Fisher Relation

### v1

@@ Abstract
The baryonic Tully Fisher relationship (bTFR) provides an empirical connection between baryonic mass and dynamical mass (measured by the maximum rotation velocity) for galaxies. Due to the impact of baryonic feedback in the shallower potential wells of dwarf galaxies, the bTFR is predicted to turn down at low masses from the extrapolated power-law relation at high masses. The low-mass end of the bTFR is poorly constrained due to small samples and difficulty in connecting the galaxy’s gas kinematics to its dark matter halo. Simulations can help us understand this connection and interpret observations. We measure the bTFR with 66 dwarf galaxies from the Marvel-ous and Marvelous Massive Dwarfs hydrodynamic simulations. Our sample has M M⊙, and is mostly gas dominated. We compare five velocity methods: V (spatially resolved mass-enclosed), V (spatially resolved midplane gravitational potential), and unresolved H i linewidths at different percentages of the peak flux (W, W, and W). We find an intrinsic turndown in the bTFR for maximum halo speeds km s-1 (or total baryonic mass, M M⊙). We find that observing H i in lower-mass galaxies to the conventional surface density limit of 1M⊙pc-2 is not enough to detect a turndown in the bTFR; none of the H i velocity methods (spatially resolved or unresolved) recover the turndown, and we find bTFR slopes consistent with observations of higher-mass galaxies. However, we predict that the turndown can be recovered by resolved rotation curves if the H i limit is M⊙ pc-2, which is within the sensitivity of current H i surveys like FEASTS and MHONGOOSE.
@@
Rotation curves provided some of the earliest evidence for dark matter halos (Bosma, 1978; Rubin et al., 1980) since rotation velocities trace the dynamical mass. The Tully Fisher Relation (TFR) (Tully & Fisher, 1977) is a measurement of galaxy stellar mass (or luminosity) as a function of rotation velocity, and thus provides an empirical constraint between a galaxy’s stellar mass and underlying dark matter halo mass. Similar to the stellar-mass halo-mass (SMHM) relation (e.g., Wechsler & Tinker, 2018; Munshi et al., 2021), the TFR helps us understand the galaxy-halo connection and galaxy evolution. Both the TFR and SMHM relation can tell us what fraction of the cosmic baryon abundance has cooled into halos and formed stars. The baryonic Tully Fisher Relation (bTFR) instead connects baryonic mass (M, total cold gas + stellar mass) to halo mass (McGaugh et al., 2000). This relation is especially useful to study lower-mass galaxies, in which the baryonic content is dominated by cold gas ( K) in the atomic hydrogen (HI) phase. Additionally, H i gas extends farther out into the halo than stars, and therefore gives us a better tracer of the underlying halo velocity to be used in the bTFR.
For galaxies with M M⊙, baryonic mass and the rotation velocity of gas follow a power-law with a bTFR slope of (e.g., Lelli et al., 2016b; Bradford et al., 2016; Lelli et al., 2019). The bTFR slope demonstrates a similar scaling between the baryonic disks and dark matter halos of galaxies, despite the many processes which redistribute baryons and dark matter (e.g., Pontzen & Governato, 2012; Di Cintio et al., 2014). Scatter in the bTFR tells us how much this scaling differs from galaxy-to-galaxy. Intrinsic scatter in the bTFR is dex, and is mostly due to scatter in the mass-concentration relation between halos (Dutton, 2012).
@@
Measuring the bTFR at lower masses is difficult due to limited sample sizes of low-surface brightness galaxies and difficulty in matching gas to halo kinematics. Only a handful of resolved rotation curves have been measured in dwarf galaxies (e.g., Oh et al., 2015). Although H i gas extends farther out in radius than stars (e.g., Krumm & Burstein, 1984; Hunter et al., 2016), the H i may not extend far enough to measure the rotation curve at the flat part. Many measurements to date extend out to a radius where the H i surface density drops below 1 M⊙ pc-2 (e.g., Broeils & Rhee, 1997; Wang et al., 2016). At this limit, some observed rotation curves are ‘still rising’, suggesting that the maximum rotation velocity is not being directly measured. McQuinn et al. (2022) constrained the bTFR with 25 galaxies at M M⊙. They used the outermost H i velocities, and found no turndown in the bTFR. However, using an adopted mass model, they were able to extrapolate the maximum rotation velocities that best matched each of their outermost rotation velocities. Assuming a cored Einasto dark matter density profile, the extrapolated velocities yielded a bTFR turndown at km s-1 and baryon fractions of 1-10% of the cosmic value.
@@
Current and upcoming radio surveys such as Apertif (van Cappellen et al., 2022; Adams et al., 2022); FEASTS (FAST Extended Atlas of Selected Targets Survey, Wang et al., 2025); MHONGOOSE (MeerKAT H i Observations of Nearby Galactic Objects: Observing Southern Emitters, de Blok et al., 2024); MIGHTEE-H i (MeerKAT International GigaHertz Tiered Extragalactice Exploration survey - HI, Maddox et al., 2021); and WALLABY (Widefield ASKAP L-band Legacy All-sky Blind surveY, Koribalski et al., 2020; Spekkens et al., 2020; Deg et al., 2024) are sensitive down to rotation speeds of km s-1, below where the bTFR turndown is expected. However, these resolved rotation curves will still be limited to the nearest dwarf galaxies. The time required to obtain large sample sizes will still be prohibitive. For representative samples, spatially unresolved H i velocities are thought to be necessary, since they require less integration time and can be measured by both radio interferometers and single-dish telescopes. Studies with unresolved H i velocities (e.g., Klypin et al., 2015; Lelli et al., 2016b; Bradford et al., 2016) can have about two orders of magnitude more galaxies than spatially resolved observations (e.g., Klypin et al., 2015; Papastergis & Shankar, 2016; McQuinn et al., 2022). Sardone et al. (2024) demonstrated that spatially unresolved, high spectral resolution observations are needed to detect the most diffuse gas in the outer wings of the H i profile, and this can reduce bTFR scatter at lower masses.
@@
It is well established that different velocity tracers of the maximum halo speed (V) exhibit different bTFR slopes. This has been studied in observations (e.g., Verheijen, 2001; Bradford et al., 2016; Lelli et al., 2019) and simulations (e.g., Brook et al., 2016; Brooks et al., 2017; Sardone et al., 2024). For spatially unresolved H i velocities, linewidths at 50% (W50) or 20% (W20) of the peak flux emission are reported as proxies for 2V. Higher-mass galaxies have flat rotation curves, and therefore more material with high column densities moving at the same velocity V. This high-velocity gas leads to a ‘double-horned’ spectral feature that makes W a better proxy to 2V. In the case of lower-mass galaxies, where H i rotation curves are still rising and the contribution from non-circular motions is greater, the H i flux profile is more Gaussian, and W50 underestimates 2V (e.g., Brook et al., 2016; Yaryura et al., 2016). W20 can provide a more accurate tracer of 2V than W50 (Macciò et al., 2016; Brook & Shankar, 2016; Brooks et al., 2017; Dutton et al., 2019; Sardone et al., 2024), though still deviates at lower masses. Although W10 (linewidth at 10% of the peak H i flux) is the largest linewidth, and therefore closest to 2Vmax for dwarf galaxies, it is also susceptible to noise from outer velocity channels.
@@
We utilize the largest suite of simulated dwarf galaxies, with realistic H i content, to understand how the bTFR turndown might manifest in observations. In this work, we analyze our sample of 66 dwarf galaxies from the Marvel-ous and Marvelous Massive Dwarfs hydrodynamic simulation suites. Our sample consists of isolated galaxies with M M⊙, and M M⊙. We focus on understanding why H i velocities (spatially resolved or unresolved) do not trace the halo’s V. We also make predictions about whether the halo’s V is recoverable with deeper H i observations. In Section 2, we describe the properties and physics implemented in our simulations, and define the velocities and masses used in our study. Section 3 presents our results and demonstrates how velocity definitions and sensitivity limits affect our ability to measure a bTFR turndown. Section 4 discusses how our bTFR results compare to observations and upcoming H i surveys. Section 5 summarizes our results.
@@ 2 Simulations & Methods
Our simulations use the code ChaNGa (Menon et al., 2015) for the SPH + N-body processes. The hydrodynamics are from GASOLINE2 (Wadsley et al., 2004, 2017), while ChaNGa has a faster gravity solver. ChaNGa scales well to simulations with a large number of particles due to its implementation in Charm++ (Kalé & Krishnan, 1993), which handles dynamic load balancing and communication for thousands of computing cores.
@@
where is the gas particle mass, is the initial star particle mass, is the star formation efficiency parameter (set to 0.1), is the H2 fraction of the gas particle, is the time step for star formation in the simulation (1 Myr), and is the local dynamical time. Star formation is limited to cold gas particles ( K), and a number density cm-3 is required. Since we also require the presence of H2, star formation usually occurs in much denser gas, so the actual density for star formation is generally cm-3. In both Marvel and Massive Dwarfs, each star particle follows the initial mass function from Kroupa (2001).
Marvel uses blastwave feedback (Stinson et al., 2006), while Massive Dwarfs uses superbubble feedback (Keller et al., 2014). Blastwave feedback injects metals, thermal energy, and mass into gas that is within the blastwave radius of the supernova (SN), as calculated by the blast wave solution for a SN remnant in Ostriker & McKee (1988). The deposited energy is erg per SN. To prevent numerical overcooling, blastwave feedback relies on turning off radiative cooling in the surrounding gas, to limit thermal energy loss in supernovae for a time corresponding to the end of the snowplow phase. Superbubble feedback does not disable radiative cooling, but instead puts feedback-heated particles into a temporary two-phase state, with each phase cooling with a separate temperature and density. These two states are kept in pressure equilibrium, and correspond roughly to the cold swept-up ISM and the hot SN-heated interior. Mass flows from the cold phase to the hot via thermal evaporation (Cowie & McKee, 1977) until the entire cold phase has been evaporated. This evaporation process can continue by stochastically evaporating cool neighbors of hot, SN-heated gas particles. Within resolved hot bubbles, temperatures gradients are smoothed by thermal conduction. The superbubble model is better able to capture the effects of multiple SNe occurring over a range of timescales, drives stronger outflows, and produces a more realistic gas phase diagram than the earlier blastwave model. For superbubble feedback, energy is deposited at a rate of erg per SN. Azartash-Namin et al. (2024) ran one set of the Marvel simulations with both feedback models, and demonstrated that they produce very similar dwarf galaxies – reproducing their expected stellar mass, size, metallicity, and bursty star formation.
@@ 2.1 Sample Selection
The faintest galaxies in most bTFR studies have stellar masses on the order of M⊙ (e.g., Bradford et al., 2015; Lelli et al., 2016a). In the case of McQuinn et al. (2022), a few galaxies have stellar masses M⊙. To better understand the bTFR turndown, our sample includes dwarf galaxies with M M⊙. We also require that each galaxy has M M⊙ in order to have enough gas for observable rotation speeds. We select isolated, central galaxies to focus on the connection between gas and dark matter kinematics without significant complications from galaxy-galaxy interactions. Bradford et al. (2016) required that any galaxy in their bTFR sample with M M⊙ be isolated, with the 2D projected distance to the nearest most massive host < 7 Mpc. All of our simulated galaxies have M M⊙, and we consider one to be ‘isolated’ if there are no galaxies more massive than the host within its virial radius. Our sample includes 66 dwarf galaxies, in which 39 are from the Massive Dwarfs suite and 27 are from the Marvel suite. Our galaxy mass ranges are (M⋆[M⊙]) = , (M[M⊙])=, and (M200[M⊙]) = . We define the total baryonic mass as M MMHI, where the factor of 1.4 accounts for helium (e.g., Arnett, 1996; Bradford et al., 2016) and metals in the gas phase. Our galaxy sample consists of 4 stellar-dominated galaxies and 62 gas-dominated galaxies, as shown in Figure 1. Based on classification using the moment of inertia tensor, we find that 34 out of 66 of our galaxies have distinct gaseous disks, while the remaining galaxies – mostly with M M⊙ – are more irregular (Keith et al., 2025).

### v2

@@ Abstract
The baryonic Tully Fisher relation (bTFR) provides an empirical connection between baryonic mass and dynamical mass (measured by the maximum rotation velocity) for galaxies. Due to the impact of baryonic feedback in the shallower potential wells of dwarf galaxies, the bTFR is predicted to turn down at low masses from the extrapolated power-law relation at high masses. The low-mass end of the bTFR is poorly constrained due to small samples and difficulty in connecting the galaxy’s gas kinematics to its dark matter halo. Simulations can help us understand this connection and interpret observations. We measure the bTFR with 75 dwarf galaxies from the Marvel-ous and Marvelous Massive Dwarfs hydrodynamic simulations. Our sample has M M⊙, and is mostly gas dominated. We compare five velocity methods: V (spatially resolved mass-enclosed), V (spatially resolved midplane gravitational potential), and unresolved H i linewidths at different percentages of the peak flux (W, W, and W). We find an intrinsic turndown in the bTFR for maximum halo speeds km s-1, or total baryonic mass M M⊙. We find that observing H i in lower-mass galaxies to the conventional surface density limit of 1 M⊙ pc-2 is not enough to detect a turndown in the bTFR; none of the H i velocity methods, spatially resolved or unresolved, recover the turndown, and we find bTFR slopes consistent with observations of higher-mass galaxies. However, we predict that the turndown can be recovered by resolved rotation curves if the H i limit is M⊙ pc-2, which is within the sensitivity of current H i surveys like FEASTS and MHONGOOSE.
@@
Rotation curves provided some of the earliest evidence for dark matter halos (Bosma, 1978; Rubin et al., 1980) since rotation velocities trace the dynamical mass. The Tully Fisher Relation (TFR) (Tully & Fisher, 1977) is a measurement of galaxy stellar mass (or luminosity) as a function of rotation velocity, and thus provides an empirical constraint between a galaxy’s stellar mass and underlying dark matter halo mass. Similar to the stellar-mass halo-mass (SMHM) relation (e.g., Wechsler & Tinker, 2018; Munshi et al., 2021), the TFR helps us understand the galaxy-halo connection and galaxy evolution. Both the TFR and SMHM relation can tell us what fraction of the cosmic baryon abundance has cooled into halos and formed stars. The baryonic Tully Fisher Relation (bTFR) instead connects baryonic mass (M, total cold gas + stellar mass) to halo mass (McGaugh et al., 2000). This relation is especially useful to study lower-mass galaxies, in which the baryonic content is dominated by cold gas (T K) in the atomic hydrogen (H i ) phase.
For galaxies with M M⊙, baryonic mass and the rotation velocity of gas follow a power-law with a bTFR slope of (e.g., Stark et al., 2009; Gurovich et al., 2010; Lelli et al., 2016b; Bradford et al., 2016; Lelli et al., 2019). The bTFR slope demonstrates a similar scaling between the baryonic disks and dark matter halos of galaxies, despite the many processes which redistribute baryons and dark matter (e.g., Pontzen & Governato, 2012; Di Cintio et al., 2014). The value of the slope has also been proposed to test different models of gravity, where a slope is consistent with modified Newtonian dynamics (Milgrom, 1983; McGaugh, 2012). Scatter in the bTFR tells us how much this scaling differs from galaxy-to-galaxy. Intrinsic scatter in the bTFR is dex, and is mostly due to scatter in the mass-concentration relation between halos (Dutton, 2012).
@@
Measuring the bTFR at lower masses is difficult due to limited sample sizes of low-surface brightness galaxies and difficulty in matching gas to halo kinematics. Only a handful of resolved rotation curves have been measured in dwarf galaxies (e.g., Oh et al., 2015). Although H i gas extends farther out in radius than stars (e.g., Krumm & Burstein, 1984; Broeils & Rhee, 1997; Begum et al., 2008; Hunter et al., 2012, 2024), the H i may not extend far out enough to where the galaxy rotates at its maximum speed. Many measurements to date extend out to a radius where the H i surface density drops below 1 M⊙ pc-2 (e.g., Broeils & Rhee, 1997; Wang et al., 2016). At this limit, some observed rotation curves are ‘still rising’, suggesting that the maximum rotation velocity is not being directly measured. McQuinn et al. (2022) constrained the bTFR with 25 galaxies at M M⊙. They used the outermost H i velocities, and found no turndown in the bTFR. However, using an adopted mass model, they were able to extrapolate the maximum rotation velocities that best matched each of their outermost rotation velocities. Assuming a cored Einasto dark matter density profile, the extrapolated velocities yielded a bTFR turndown at km s-1 and baryon fractions of 1-10% of the cosmic value.
@@
Current and upcoming radio surveys such as Apertif (van Cappellen et al., 2022; Adams et al., 2022); FEASTS (FAST Extended Atlas of Selected Targets Survey, Wang et al., 2025); MHONGOOSE (MeerKAT H i Observations of Nearby Galactic Objects: Observing Southern Emitters, de Blok et al., 2024); MIGHTEE-H i (MeerKAT International GigaHertz Tiered Extragalactice Exploration survey - HI, Maddox et al., 2021); and WALLABY (Widefield ASKAP L-band Legacy All-sky Blind surveY, Koribalski et al., 2020; Spekkens et al., 2020; Deg et al., 2024) are sensitive down to expected H i masses of galaxies with rotation speeds of km s-1, below where the bTFR turndown is expected. However, these resolved rotation curves will still be limited to the nearest dwarf galaxies. The time required to obtain large sample sizes will still be prohibitive. For representative samples, spatially unresolved H i velocities are thought to be necessary, since they require less integration time and can be measured by both radio interferometers and single-dish telescopes. Studies with unresolved H i velocities (e.g., Klypin et al., 2015; Lelli et al., 2016b; Bradford et al., 2016) can have about two orders of magnitude more galaxies than spatially resolved observations (e.g., Klypin et al., 2015; Papastergis & Shankar, 2016; McQuinn et al., 2022). Sardone et al. (2024) demonstrated that spatially unresolved, high spectral resolution observations are needed to detect the most diffuse gas in the outer wings of the H i profile, and this can reduce bTFR scatter at lower masses.
@@
It is well established that different velocity tracers of the maximum halo speed (V) exhibit different bTFR slopes. This has been studied in observations (e.g., Verheijen, 2001; Bradford et al., 2016; Lelli et al., 2019) and simulations (e.g., Brook et al., 2016; Brooks et al., 2017; Glowacki et al., 2020; Sardone et al., 2024). For spatially unresolved H i velocities, linewidths at 50% (W50) or 20% (W20) of the peak flux emission are reported as proxies for 2V. Higher-mass galaxies have flat rotation curves, and therefore more material with high column densities moving at the same velocity V. This high-velocity gas leads to a ‘double-horned’ spectral feature that makes W a better proxy to 2V. In the case of lower-mass galaxies, where H i rotation curves are still rising and the contribution from non-circular motions is greater (El-Badry et al., 2018), the H i flux profile is more Gaussian, and W50 underestimates 2V (e.g., Brook et al., 2016; Yaryura et al., 2016). W20 can provide a more accurate tracer of 2V than W50 (Macciò et al., 2016; Brook & Shankar, 2016; Brooks et al., 2017; Dutton et al., 2019; Sardone et al., 2024), though still deviates at lower masses. Although W10 (linewidth at 10% of the peak H i flux) is the largest linewidth, and therefore closest to 2Vmax for dwarf galaxies, it is also susceptible to noise from outer velocity channels.
@@
We utilize the largest suite of simulated dwarf galaxies, with realistic H i content, to understand how the bTFR turndown might manifest in observations. In this work, we analyze our sample of 75 dwarf galaxies from the Marvel-ous and Marvelous Massive Dwarfs hydrodynamic simulation suites. Our sample consists of isolated galaxies with M M⊙, and M M⊙. We focus on understanding why H i velocities (spatially resolved or unresolved) do not trace the halo’s V. We also make predictions about whether the halo’s V is recoverable with deeper H i observations. In Section 2, we describe the properties and physics implemented in our simulations, and define the velocities and masses used in our study. Section 3 presents our results and demonstrates how velocity definitions and sensitivity limits affect our ability to measure a bTFR turndown. Section 4 discusses how our bTFR results compare to observations and upcoming H i surveys. Section 5 summarizes our results.
@@ 2 Simulations & Methods
Our simulations use the code ChaNGa (Menon et al., 2015) for the SPH + N-body processes. The hydrodynamics are from GASOLINE2 (Wadsley et al., 2004, 2017), while ChaNGa has a faster gravity solver. ChaNGa scales well to simulations with a large number of particles due to its implementation in Charm++ (Kale & Krishnan, 1993), which handles dynamic load balancing and communication for thousands of computing cores.
@@
where is the gas particle mass, is the initial star particle mass, is the star formation efficiency parameter (set to 0.1), is the H2 fraction of the gas particle, is the time step for star formation in the simulation (1 Myr), and is the local dynamical time. Star formation is limited to cold gas particles ( K), and a number density cm-3 is required. Since we also require the presence of H2, star formation usually occurs in much denser gas, so the actual density for star formation is generally cm-3. Our star formation prescription from Christensen et al. (2012) reproduces the Kennicut-Schmidt relation (Schmidt, 1959; Kennicutt, 1998) – e.g., see right panel of Figure 10 in Christensen et al. (2014). In both Marvel and Massive Dwarfs, each star particle follows the initial mass function from Kroupa (2001).
@@
Marvel uses blastwave feedback (Stinson et al., 2006), while Massive Dwarfs uses superbubble feedback (Keller et al., 2014). Blastwave feedback injects metals, thermal energy, and mass into gas that is within the blastwave radius of the supernova (SN), as calculated by the blastwave solution for a SN remnant in Ostriker & McKee (1988). The deposited energy is erg per SN. To prevent numerical overcooling, blastwave feedback relies on turning off radiative cooling in the surrounding gas, to limit thermal energy loss in supernovae for a time corresponding to the end of the snowplow phase. Superbubble feedback does not disable radiative cooling, but instead puts feedback-heated particles into a temporary two-phase state, with each phase cooling with a separate temperature and density. These two states are kept in pressure equilibrium, and correspond roughly to the cold swept-up ISM and the hot SN-heated interior. Mass flows from the cold phase to the hot via thermal evaporation (Cowie & McKee, 1977) until the entire cold phase has been evaporated. This evaporation process can continue by stochastically evaporating cool neighbors of hot, SN-heated gas particles. Within resolved hot bubbles, temperatures gradients are smoothed by thermal conduction. The superbubble model is better able to capture the effects of multiple SNe occurring over a range of timescales, drives stronger outflows, and produces a more realistic gas phase diagram than the earlier blastwave model. For superbubble feedback, energy is deposited at a rate of erg per SN. Superbubble feedback is thought to have stronger outflows that expel low angular momentum gas and regulate star formation (Mina et al., 2021), yet Piacitelli et al. (2025) find no difference in the circumgalactic medium in a larger sample of galaxies combining the two feedback models.
Although these simulations have different physics, we use galaxies from both in order to span a larger mass range. The Marvel-ous Dwarf volumes are dominated by lower-mass dwarfs, while the Massive Dwarfs simulations are (by definition) focused on higher stellar mass dwarfs of M M⊙, and any additional field dwarf galaxies contained in the zoom region. We expect properties like mass, size, and rotation velocity to generally be consistent across simulations. Azartash-Namin et al. (2024) ran one set of the Marvel simulations with blastwave and superbubble feedback models, and demonstrated that both feedback models reproduced the observed trends for Local Volume dwarf galaxies in terms of their stellar mass, size, and metallicity. Galaxies with superbubble feedback exhibited higher halo masses, though the variance between feedback models was within the observational error of the SMHM relation. This result suggests that galaxies from the Massive Dwarfs simulation may have slightly higher rotation velocities in a given mass range (compared to galaxies from Marvel), but we expect the difference to be small.
Piacitelli et al. (2025) also used galaxies from Marvel and Massive Dwarfs (with blastwave and superbubble feedback, respectively) and found consistent trends between simulations for specific star formation rate versus stellar mass, the luminosity-metallicity relation, and column density profiles for gas in the circumgalactic medium. Meanwhile, Munshi et al. (2021) showed that the slightly different resolutions used in these simulation sets has no impact on the resulting SMHM relation. In summary, there is no conclusive evidence that the differing feedback models or resolutions yield significantly different galaxy properties for the two suites. Given these results, we combine them here for a sample that spans as large a mass range as possible.
@@ 2.1 Sample Selection
The faintest galaxies in most bTFR studies have stellar masses on the order of M⊙ (e.g., Bradford et al., 2015; Lelli et al., 2016a). In the case of McQuinn et al. (2022), a few galaxies have stellar masses M⊙. To better understand the bTFR turndown, our sample includes dwarf galaxies with M M⊙. We also require that each galaxy has M M⊙ in order to have enough gas for observable rotation speeds. We select isolated field galaxies to focus on the connection between gas and dark matter kinematics without significant complications from galaxy-galaxy interactions. Bradford et al. (2016) required that any galaxy in their bTFR sample with M M⊙ be isolated, with the 2D projected distance to the nearest most massive host > 1.5 Mpc. All of our simulated galaxies have M M⊙, and are only included if there are no other galaxies of equal or greater mass within the virial radius. Our sample includes 75 dwarf galaxies, of which 48 are from the Massive Dwarfs suite and 27 are from the Marvel suite. Our galaxy mass ranges are (M⋆[M⊙]) = , (M[M⊙])=, and (M200[M⊙]) = . Mass and size distributions for our sample are shown in Figure 1. We define the total baryonic mass as M MMHI, where the factor of 1.4 accounts for helium (e.g., Arnett, 1996; Bradford et al., 2016) and metals in the gas phase. Our galaxy sample consists of 5 stellar-dominated galaxies and 70 gas-dominated galaxies, as shown in Figure 1. Using these simulations, Keith et al. (2025) found that galaxies with M M⊙ are typically more irregular while those at higher masses have disks. They determined this by using the principle axis ratios of the 3D ellipsoid that best fit the shape of the star particles, and the principle axes were calculated from the stellar moment of inertia tensor.